# PhyMPGN: Physics-encoded Message Passing Graph Network for spatiotemporal PDE systems

### v1

@@ 4.3 Ablation study
Additionally, using the same datasets as before, we investigate the impact of three types of numerical integrators employed in our model on performance: the Euler forward scheme, the RK2 scheme, and the RK4 scheme. The MSEs of our models employing these three numerical integrators are presented in Table A.11, showing that the higher the accuracy of the integrator, the better the model’s prediction accuracy. However, higher-order integrators also come with greater memory and computational costs. Moreover, empirical evidence suggests that RK4 may increase the difficulty of training, including issues such as training instability. Therefore, we choose RK2 in this paper to strike a balance between computational resources and accuracy. Notably, even with the Euler scheme, PhyMPGN’s accuracy remains superior to other baseline models with a big margin.
@@ 5 Conclusion
We present a graph learning approach, namely, PhyMPGN, for prediction and simulation of spatiotemporal PDE systems on coarse unstructured meshes given small training datasets. Specifically, we develop a physics-encoded message-passing GNN model, where the temporal marching is realized via a second-order numerical integrator (e.g. Runge-Kutta scheme). The a priori physics knowledge is embedded to aid and guide the GNN learning in a physically feasible solution space with improved solution accuracy, via introducing (1) a learnable Laplace block that encodes the discrete Laplace-Beltrami operator, and (2) a novel padding strategy to encode different types of BCs. Extensive experiments demonstrate that PhyMPGN outperforms other baseline models with considerable margins, e.g., exceeding 50% gains across diverse spatiotemporal dynamics, given small and sparse training datasets. However, several challenges remain to be addressed: (1) how to effectively encode Neumann/Robin BCs in latent space; (2) extending the model to 3-dimensional scenarios; These challenges highlight key directions for our future research.
@@ Appendix C Encoding boundary conditions
Before constructing the unstructured mesh by Delaunay triangulation, we first apply the padding strategy to the discrete nodes in the physical space (i.e., $\bm{u}^{k}$), as shown in Figure A.10. We consider four types of BCs: Dirichlet, Neumann, Robin, and periodic. All nodes on the Dirichlet boundary will be padded with the specified values. For Neumann/Robin BCs, ghost nodes are created to be symmetric with the nodes nearby the boundary in the physical space, and their padded values are based on gradient information in the normal direction. For periodic BCs, the nodes near the boundary are flipped to the other side, achieving a cyclic effect in message passing (detailed formulations of various BCs are provided in Table A.3). Once the padding is completed, Delaunay triangulation is applied to construct the mesh, which serves as the graph input for the model. Apart from this, we also apply padding to the prediction (i.e., $\bm{u}^{k+1}$) of the model at each time step to ensure that it satisfies the BCs before being fed into the model for next-step prediction, as shown in Figure A.11.
@@ Appendix E Modeling PDE systems
We train and evaluate PhyMPGN’s performance on various physical systems, including Burgers’ equation, FN equation, GS equation and cylinder flow. Leveraging four NVIDIA A100 GPUs, the training process are completed under 12 hours. And the inference time are ranging from several seconds to several minutes. We also compare PhyMPGN against existing approaches, such as DeepONet, PA-DGN, MGN, MP-PDE.
@@
Additionally, using the same datasets as before, we investigate the impact of three types of numerical integrators employed in our model on performance: the Euler forward scheme, the RK2 scheme, and the RK4 scheme. The MSEs of our models employing these three numerical integrators are presented in Table A.11.

### v2

@@ 4.3 Ablation study
Additionally, we investigate the impact of three types of numerical integrators employed in our model on performance: Euler forward, RK2, and RK4 scheme. The MSEs of these models are presented in Table A.11, showing that the higher the order of the integrator, the better the model’s prediction accuracy. However, higher-order integrators also come with greater memory and computational costs. Moreover, empirical evidence suggests that RK4 may increase the difficulty of training, including issues such as training instability. Therefore, we choose RK2 to strike a balance between computational resources and accuracy. Notably, even with Euler scheme, PhyMPGN’s accuracy remains superior to baseline models with a big margin. We also explore the impact of the time segment size $M$ on results and computational requirements. Details are presented in Appendix E.
@@ 5 Conclusion
We present a graph learning approach, namely, PhyMPGN, for predicting spatiotemporal PDE systems on coarse unstructured meshes given small training datasets. Specifically, we develop a physics-encoded message-passing GNN model, where the temporal marching is realized via a second-order numerical integrator (e.g. Runge-Kutta scheme). The a priori physics knowledge is embedded to aid and guide the GNN learning in a physically feasible solution space with improved solution accuracy, via introducing (1) a learnable Laplace block that encodes the discrete Laplace-Beltrami operator, and (2) a novel padding strategy to encode different types of BCs. Extensive experiments demonstrate that PhyMPGN outperforms other baseline models with considerable margins, e.g., exceeding 50% gains across diverse spatiotemporal dynamics, given small and sparse training datasets. However, several challenges remain to be addressed: (1) how to effectively encode Neumann/Robin BCs in latent space; (2) the approach cannot be applied to other nonlinear or high-order diffusion terms (e.g., $\bm{\nabla}^{4}$); (3) extending the model to 3-dimensional scenarios. These challenges highlight key directions for our future research.
Acknowledgments
The work is supported by the National Natural Science Foundation of China (No. 62276269 and No. 92270118), the Beijing Natural Science Foundation (No. 1232009), and the Fundamental Research Funds for the Central Universities (No. 202230265).
APPENDIX
@@ Appendix C Encoding boundary conditions
Before constructing the unstructured mesh by Delaunay triangulation, we first apply the padding strategy to the discrete nodes in the physical space (i.e., $\bm{u}^{k}$), as shown in Figure A.2. We consider four types of BCs: Dirichlet, Neumann, Robin, and periodic. All nodes on the Dirichlet boundary will be directly assigned with specified values. For Neumann/Robin BCs, ghost nodes are created symmetrically with respect to the nodes near the boundary (along the normal direction of the boundary) in the physical space, and their padded values depends on derivatives in the normal direction. The goal is to ensure the true nodes, the boundary, and the ghost nodes satisfy the BCs in a manner similar to the central difference method. For periodic BCs, the nodes near the boundary $\Gamma_{p_{1}}$ are flipped and placed near the corresponding boundary $\Gamma_{p_{2}}$, achieving a cyclic effect in message passing (detailed formulations of various BCs are provided in Table A.2). Once the padding is completed, Delaunay triangulation is applied to construct the mesh, which serves as the graph input for the model. Apart from this, we also apply padding to the prediction (i.e., $\bm{u}^{k+1}$) of the model at each time step to ensure that it satisfies the BCs before being fed into the model for next-step prediction, as shown in Figure A.3.
@@
PENN and our method both encode BCs. However, there are several key differences between the two approaches, namely,
  • Methods: PENN designs the special neural layers and modules for the model to satisfy BCs while we develop a padding strategy directly applied to the features.
  • Types of BCs: PENN proposes DirichletLayer and pseudoinverse decoder for Dirichlet BCs, and NeumannIsoGCN for Neumann BCs. In contrast, our padding strategy can be applied to four types of BCs (Dirichlet, Neumman, Robin, periodic), which can be seamlessly integrated into any GNN model and thus has a better generality.
  • Accuracy: Both PENN and our padding strategy are designed to completely satisfy Dirichlet BCs without any error. However, neither method can enforce other BCs without error.
  • Efficiency: During training, the additional cost for PENN to satisfy BCs involves constructing the pseudo-inverse decoder after the parameters are updated. The efficiency of this process depends on the number of parameters in the encoder. In contrast, the cost of our padding strategy involves mere copying the corresponding nodes in the domain during both training and testing, which is small.
To quantify the time overhead of our padding strategy, taking Burgers’ trajectories with 1600 time steps and 503 nodes in the domain as an example, we compare the inference time of our models with and without the padding strategy. As shown in Table A.1, the results demonstrate that the overhead introduced by the padding strategy is small.
@@
Table A.1: The inference time of models with and without the padding strategy.
Inference time (s)
PhyMPGN (padding) 10.71
PhyMPGN (without padding) 10.02
@@ Appendix E Modeling PDE systems
We train and evaluate PhyMPGN’s performance on various physical systems, including Burgers’ equation, FN equation, GS equation and cylinder flow. Leveraging four NVIDIA RTX 4090 GPUs, the training process is completed within a range of 4 to 15 hours. And the inference time is under one minute on a single GPU. We also compare the performance of PhyMPGN against existing approaches, such as DeepONet, PA-DGN, MGN, MP-PDE.
@@
Table A.10: The inference time (seconds) and MSE of four methods for Burgers’ equation.
Time per trajectory Total time (10 trajectories) MSE
MGN 4.01 5.19 9.19e-4
MP-PDE 0.33 1.00 5.88e-3
COMSOL 19.5 195 4.91e-5
PhyMPGN 10.71 13.64 2.99e-4
@@
Computational efficiency:
The computational cost of the Laplace block and the padding strategy is minimal. The computation of Mesh Laplace in Laplace block includes only matrix multiplication $L_{w}\bm{f}$, where $\bm{f}$ are the node features and $L_{w}$ is a weight matrix related to mesh geometric that can be calculated offline. The cost of the padding strategy involves merely copying the corresponding nodes in the domain and is small.
For a comparative analysis of computational efficiency, we select MGN and MP-PDE as baselines from neural-based methods, and COMSOL as a baseline from classical numerical methods. Taking Burgers’ trajectories with 1600 time steps and 503 nodes in the domain as an example, we compare the inference times between these three baselines and our model (PhyMPGN). The inference time and MSE of these four methods are shown in Table A.10. All evalutions are conducted on a single NVIDIA RTX 4090 GPU and an i9-13900 CPU. Additionally, the ground truth is generated with high resolution in both time and space using COMSOL with the implicit scheme.
MGN predicts one step given one step as input, similar to PhyMPGN. However, due to the additional Laplace block, padding strategy, and RK2 scheme, PhyMPGN has more inference time than MGN. On the other hand, MP-PDE predicts $T_{w}$ steps (with $T_{w}=20$) in a single forward pass using $T_{w}$ steps as input, resulting in the least inference time among neural-based models. Overall, these neural-based models can greatly improve efficiency by simultaneously inferring multiple trajectories. The results from COMSOL, a multiphysics simulation software based on advanced numerical methods, presented in the comparison, are obtained using an implicit scheme.
In summary, PhyMPGN achieves the lowest error among neural-based models with a reasonable time overhead as a trade-off, while COMSOL delivers the least error among these four methods but at the cost of significantly higher time overhead. Notably, to simulate the PDE system, COMSOL requires the complete governing PDE formulas, whereas the three neural-based methods do not, as they learn the dynamics from data.
@@
Investigation on numerical integrators:
Additionally, using the same datasets as before, we investigate the impact of three types of numerical integrators employed in our model on performance: the Euler forward scheme, the RK2 scheme, and the RK4 scheme. The MSEs of our models employing these three numerical integrators are presented in Table A.11. We trained our model for 1600 epochs using four NVIDIA RTX 4090 GPUs and performed inference for 2000 time steps per trajectory on a single GPU. Their training and inference time are listed in Table A.12. All inference times are under one minute. Thanks to the parallel capabilities of GPUs, the inference speed can be further accelerated by processing batches simultaneously.
Table A.12: The training and inference time of our models employing different types of numerical integrators for cylinder flow problems.
Training time Inference time per trajectory Inference time (5 trajectories)
PhyMPGN (Euler) 5.2 h 8.98 s 13.47 s
PhyMPGN (RK2) 14.4 h 17.57 s 27.08 s
PhyMPGN (RK4) 39.3 h 34.99 s 53.49 s
Table A.13: The MSE and the training time with different segment sizes $M$.
$M$ $Re=[200,280,360]$ $Re=[440,480,520]$ $Re=[600,800,1000]$ Training time
10 2.43e-3 1.48e-1 2.42e-1 8.3 h
15 9.34e-4 3.05e-2 1.34e-1 15.0 h
20 2.14e-4 3.56e-2 1.25e-1 14.4 h
The impact of the time segment size $M$:
Many previous methods (Seo et al., 2020; HAN et al., 2022; Geneva & Zabaras, 2022; Ren et al., 2022; Rao et al., 2023) unroll models to predict multiple steps within a short time segment then backpropagate to learn the dynamics of physics systems. Intuitively, longer time segments can be advantageous for capturing long-term dependencies, thereby improving the model’s long-term prediction capabilities. However, excessively long segments may introduce challenges during training, such as increased memory consumption and difficulties in achieving convergence. To investigate the impact of the segment size $M$ on the results and computational requirements, we conduct experiments using the cylinder flow datasets in the ablation study. We train our model with segment sizes $M=10,15,20$, and the MSEs on the testing sets and the training time are presented in Table A.13, which shows that the bigger the segment size $M$, the better the model’s prediction accuracy. Additionally, since we need to adjust the batch size to fit the GPU memory as the segment size $M$ varies, the training time does not scale linearly with $M$. Specifically, the batch sizes for $M=15$ and $M=20$ are the same, while the batch size for $M=10$ is double that of the other two configurations. Therefore, the training time differences are also reasonable.